### v1

# Robust Learning in Bayesian Parallel Branching Graph Neural Networks: The Narrow Width Limit

Abstract
The infinite width limit of random neural networks is known to result in Neural Networks as Gaussian Process (NNGP) (Lee et al. (2018)), characterized by task-independent kernels. It is widely accepted that larger network widths contribute to improved generalization (Park et al. (2019)). However, this work challenges this notion by investigating the narrow width limit of the Bayesian Parallel Branching Graph Neural Network (BPB-GNN), an architecture that resembles residual networks. We demonstrate that when the width of a BPB-GNN is significantly smaller compared to the number of training examples, each branch exhibits more robust learning due to a symmetry breaking of branches in kernel renormalization. Surprisingly, the performance of a BPB-GNN in the narrow width limit is generally superior or comparable to that achieved in the wide width limit in bias-limited scenarios. Furthermore, the readout norms of each branch in the narrow width limit are mostly independent of the architectural hyperparameters but generally reflective of the nature of the data. Our results characterize a newly defined narrow-width regime for parallel branching networks in general.
1 Introduction
The study of neural network architectures has seen substantial growth, particularly in understanding how network width impacts learning and generalization. It is generally believed that wider networks generally perform better (Allen-Zhu et al. (2019); Jacot et al. (2018); Gao et al. (2024)). However, this work challenges the prevailing assumption by exploring the narrow width limit of Bayesian Parallel Branching Graph Neural Networks (BPB-GNNs), an architecture inspired by residual GCN networks (Chen et al. (2020a, 2022)). We show theoretically and empirically that narrow-width networks can perform better than their wider counterparts due to a symmetry-breaking effect in kernel renormalization, in bias-limited scenarios. This paper presents a detailed theoretical analysis of BPB-GNNs in the narrow-width regime, highlighting realistic conditions under which these networks demonstrate robust learning and comparable generalization.
Contributions
  1. 1.
    We introduce a novel yet simple GCN architecture with parallel independent branches, and derive the exact generalization error for node regression in the statistical limit as the sample size $P\to\infty$ and network width $N\to\infty$, with their ratio a finite number $\alpha=P/N$, in the over-parametrized regime.
  2. 2.
    We show that in the Bayesian setting, the bias will decrease and saturate at narrow hidden layer width, a surprising phenomenon due to kernel renormalization. We demonstrate that this can be understood as a robust learning effect of each branch in the student-teacher task, where each student branch is learning the teacher’s branch.
  3. 3.
    We demonstrate this narrow-width limit in real-world dataset Cora and understand each branch’s importance as a nature of the dataset.
2 Related works
Infinitely wide neural networks:
Our work follows a long tradition of mathematical analysis of infinitely-wide neural networks (Jacot et al. (2018); Lee et al. (2018); Bahri et al. (2024)), resulting in NTK or NNGP kernels. Recently, such analysis has been extended to structured neural networks including GNNs (Du et al. (2019); Walker and Glocker (2019)). However, they do not provide analysis of feature learning where the kernel is dependent on the tasks.
Kernel renormalization and feature learning:
There has been progress on understanding simple MLPs in the feature-learning regime as the kernel shape changes with the task or time (Li and Sompolinsky (2021); Atanasov et al. (2021); Avidan et al. (2023); Wang and Jacot (2023)). We develop such understanding in Graph-based networks.
Theoretical analysis of GCN:
There is a long line of works that analyze theoretically the expressiveness (Xu et al. (2018); Geerts and Reutter (2022)) and generalization performance (Tang and Liu (2023); Garg et al. (2020); Aminian et al. (2024)) of GNN. However, it is challenging to calculate the generalization error dependence on tasks. To our knowledge, our work is first to provide a tight bound of the generalization error for GNN with residual-like structures. The closest architecture to our linear BPB-GCN is the linearly decoupled GCN proposed by Cong et al. (2021); however, the readout vector in their case has weight-sharing, which will not result in kernel renormalization for different branches.
3 BPB-GNN
We are motivated to study the parallel branching networks as they resemble residual blocks in commonly used architectures and tractable to study analytically with our Bayesian framework. Given graph $G=(A,X)$, where $A$ is the adjacency matrix and $X$ the node feature matrix, the final readout for node $\mu$ is a scalar $f^{\mu}(G;\Theta)$ which depends on the graph and network parameters $\Theta$.
3.1 Parallel branching GNN architecture
The parallel branching GNN is an ensemble of GNN branches, where each branch operates independently with no weight sharing. In this work, we focus on the simple setup of branches made of linear GCN with one hidden layer, but with different number of convolutions $A$ [^1] In this paper, the convolution operation is normalized as $A=D^{-1/2}(\hat{A}+I)D^{-1/2}$, where $\hat{A}$ is the original Adjacency matrix and $D$ is the degree matrix. We also use feature standardization after convolution for each branch to normalize the input. on the input node features (Figure 1(a)). In this way, the parallel branching GNN is analogous to GCN with residual connections, for which the final node readout can also be thought of as an ensemble of convolution layers (Veit et al. (2016)). Concretely, the overall readout $f^{\mu}(G;\Theta)$ for node $\mu$ is a sum of the branch readouts
$f^{\mu}(G;\Theta)=\sum_{l=0}^{L-1}f_{l}^{\mu}(G;\Theta_{l}=\{W^{(l)},a^{(l)}\}),$ (1)
where
$f_{l}^{\mu}(G,\Theta_{l})=\frac{1}{\sqrt{L}}\sum_{i=1}^{N}\frac{1}{\sqrt{N}}a_{i}^{(l)}\sum_{j=1}^{N_{0}}\frac{1}{\sqrt{N_{0}}}W_{ij}^{(l)}\sum_{\nu=1}^{n}(A^{l})_{\mu\nu}x_{j}^{\nu}$ (2)
In matrix notation,
$F=\sum_{l}\frac{1}{\sqrt{LNN_{0}}}A^{l}XW^{(l)}a^{(l)}$ (3)
Note that when $L=2$, the BPB-GNN reduces exactly to a 2-layer residual GCN (Chen et al. (2020b)).
(a) The parallel branching GNN architecture, with 2 branches.
(b) Student and teacher readout norms squared for wide and narrow student BPB-GNN networks.
Figure 1: Overview of the main message of the paper: narrow BPB-GNN learns robust representations for each branch at narrow width.
3.2 Bayesian node regression
We consider a Bayesian semi-supervised node regression problem, for which the posterior probability for the weight parameters is given by
$P(\Theta)=\frac{1}{Z}e^{-E(\Theta;G,Y)/T}=\frac{1}{Z}\exp({-\frac{1}{2T}\sum_{\mu=1}^{P}(f^{\mu}(G,\Theta)-y^{\mu})^{2}-\frac{1}{2\sigma^{2}}\Theta^{T}\Theta}),$ (4)
[…]
We are interested in understanding the weight and predictor statistics of each branch and how they contribute to the overall generalization performance of the network. In the following theoretical derivations, our working regime is in the overparametrizing high dimensional limit (Li and Sompolinsky (2021); Montanari and Subag (2023); Bordelon and Pehlevan (2022); Howard et al. (2024)): $P,N,N_{0}\to\infty$, $\frac{P}{N}=\alpha$ finite, and the capacity $\alpha_{0}=\frac{P}{LN_{0}}$<1. As we will show later, this limit is practically true even with $P,N$ not so large (our smallest $N$ is 4). We will also use near 0 temperature in which case the training error will be near 0 and the prior $L_{2}$ regularization has an inductive bias on the solution space that will influence the generalization properties.
3.3 Kernel renormalization and order parameters
[…]
where the effective Hamiltonian $H(W)$ in terms of the hidden layer weights for all branches is
$H(W)=\frac{1}{2\sigma^{2}}\sum_{l=0}^{L-1}\mathrm{Tr}W_{l}^{T}W_{l}+\frac{1}{2}Y^{T}(K(W)+TI)^{-1}Y+\frac{1}{2}\log\det(K(W)+TI),$ (6)
where
$K(W)=\frac{1}{L}\sum_{l}\frac{\sigma^{2}}{N}(H_{l}(W_{l})H_{l}(W_{l})^{T})|_{P}$ (7)
is the $P\times P$ kernel matrix dependent on the observed $P$ nodes with node features $H_{l}=A^{l}XW_{l}$ and we denote $|_{P}$ as the matrix restricting to the elements generated by the training nodes.
As shown in Appendix A.1, we can further integrate out the $W_{l}$’s and get the partition function $Z=\int\,Due^{-H(u)}$ described by a final effective Hamiltonian independent of weights
$H(u)=S(u)+E(u),$ (8)
where we call $S(u)$ the entropic term
$S(u)=-\sum_{l}\frac{N}{2}\log u_{l}+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}$ (9)
and $E(u)$ the energetic term
$E(u)=\frac{1}{2}Y^{T}(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)^{-1}Y+\frac{1}{2}\log\det(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI),$ (10)
where $K_{l}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{P}$ is the ($P\times P$) input node feature kernel.
Therefore, the final effective Hamiltonian has the overall kernel
$K=\sum_{l}\frac{1}{L}u_{l}K_{l},$ (11)
where $u_{l}$’s are order parameters which is the minimum of the effective Hamiltonian Eq. 8 by saddle point methods.
GP kernel vs. renormalized kernel:
Observe that as $\alpha=P/N\to 0$, the entropic term dominates, and thus $u_{l}=\sigma^{2}$ for all branches. This is the usual Gaussian Process (GP) limit, or the infinite-width limit. In this case, the kernel $K_{l}$ for each branch will not be changed by the training data and each branch has the same contribution in terms of its strength.
However, when $\alpha=P/N$ is large, there is going to be a correction to the GP prediction and $u_{l}$’s in general are going to depend on the training data; we therefore have feature learning in each branch with kernel renormalization. It turns out that $u_{l}$ is exactly the statistical average of readout norm squared (Appendix A.3)
$u_{l}=\langle\|a_{l}\|^{2}\rangle/N$ (12)
for each branch $l$; therefore, branches in general become more and more different as the hidden layer width $N$ gets smaller. We call this phenomenon symmetry breaking, which is discussed in Section 4.2.
3.4 Predictor statistics and generalization
Under the theoretical framework, we obtain analytically (Appendix A.2) the mean prediction and variance for a single test node $\nu$ as
$\langle y^{\nu}(G)\rangle=\sum_{l}\frac{1}{L}u_{l}k_{l,\nu}(K+TI)^{-1}Y$ (13)
$\langle\delta y_{\nu}(G)^{2}\rangle=K_{\nu,\nu}-\sum_{l^{\prime}}\frac{1}{L}u_{l^{\prime}}k_{l^{\prime},\nu}^{T}(K+TI)^{-1}\sum_{l^{\prime\prime}}\frac{1}{L}u_{l^{\prime\prime}}k_{l^{\prime\prime},\nu}$ (14)
Here
$k_{l}^{\nu}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{(P,\nu)}$ (15)
is the $P\times 1$ column kernel matrix for test node $\nu$ and all training nodes, and
$K^{\nu,\nu}=\frac{\sigma^{2}}{N_{0}}\sum_{l}\frac{1}{L}u_{l}[A^{l}XX^{T}A^{l}]|_{(\nu,\nu)}$ (16)
is the single matrix element for the test node. We will use this theoretical prediction to calculate the generalization performance, defined by the MSE on $t$ test nodes
$\epsilon_{g}=\langle\frac{1}{t}\sum_{\nu=1}^{t}(f^{\nu}(G)-y^{\nu})^{2}\rangle_{\Theta}=\mathrm{Bias}+\mathrm{Variance},$ (17)
where
$\mathrm{Bias}=\frac{1}{t}\sum_{\nu=1}^{t}(\langle f^{\nu}(G)\rangle_{\Theta}-y^{\nu})^{2},\mathrm{Variance}=\frac{1}{t}\sum_{\nu=1}^{t}\langle\delta f^{2}_{\nu}\rangle.$ (18)
Note that our definition of bias and variance is a statistical average over the posterior weight distribution, which is slightly different from the usual definition from GD. However, we expect that the overall generalization error to be similar to the one resulting from one random initialization, when the test sample size is large enough.
4 The narrow width limit
As we discussed briefly in 3.3, the kernel becomes highly renormalized at narrow width. In fact, in the other extreme scenario when $N/P\to 0$, the energetic term in the Hamiltonian completely dominates, and we would expect that the generalization performance saturates as the order parameters in the energetic terms become independent of width $N$. Therefore, just as infinitely wide networks correspond to the GP limit, we propose that there exists a narrow width limit when the network width is extremely small compared to the number of training samples.
4.1 Robust learning of branches: the equipartition conjecture
What happens in the narrow width limit? In the following, we demonstrate that each branch will learn robustly at narrow width.
Consider a student-teacher network setup, where the teacher network is given by
$f^{*}(G;\Theta^{*})=\sum_{l}f_{l}^{*}(G;W_{l}^{*})=\sum_{l}\frac{1}{\sqrt{NL}}\sum_{i}a_{i,l}^{*}h_{i}^{l}(G;W_{l}^{*}).$ (19)
$W_{ij}^{*}\sim\mathcal{N}(0,\sigma_{t}^{2})$ and $a_{i,l}^{*}\sim\mathcal{N}(0,\beta_{l}^{2})$, where $\beta_{l}^{2}$ is the assigned variance of the readout weight for the teacher branch $l$. Now we feed the data generated from this distribution $Y$ to the student network with the same number of branches, with student variance $\sigma^{2}$.
Consider $(YY^{T})_{\mu,\nu}=\sum_{i,j,l_{1},l_{2}}\frac{1}{NL}a^{*}_{i,l_{1}}a^{*}_{j,l_{2}}h_{i}^{l_{1}}(G)h_{j}^{l_{2}}(G)$; we conjecture that this quantity concentrates at its expectation value $\mathbf{E}_{a^{*},W^{*}}(YY^{T})=\sum_{l}\beta_{l}^{2}K_{l}/L$. Given this assumption, we can calculate the order parameters in the saddle point equation 40 (A.1) at 0 temperature. Ignoring the entropic term at narrow width, $r_{l}$ (Appendix A) becomes
$r_{l}=Y^{T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y\approx\mathrm{Tr}K^{-1}\frac{u_{l}K_{l}}{L}(K^{-1}\sum_{l}\beta_{l}^{2}\frac{K_{l}}{L})=\mathrm{tr}_{l}(K^{-1}\sum_{l}\beta_{l}^{2}\frac{K_{l}}{L})$ (20)
One solution that satisfies the saddle point equations is
$u_{l}\sigma^{2}=\beta_{l}^{2}\sigma_{t}^{2}.$ (21)
We call this the equipartition conjecture, as the mean readout squared and the variance (A.3) has to exactly balance each other. Furthermore, by Eq.12, the student readout norm at each branch on average learns the teacher readout norm when the variance of teacher network $\sigma_{t}^{2}=\sigma^{2}$ and $N\to\infty$, ie.
$\langle\|a_{\mathrm{student}}\|^{2}\rangle=\|a_{\mathrm{teacher}}\|^{2}$ (22)
Furthermore, writing $u_{l}K_{l}/L\approx Y_{l}Y^{T}$, at narrow width, the predictor statistics for each stream for the training data from the teacher becomes
$\langle f_{l}(X)\rangle=\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y\approx Y_{l},$ (23)
ie. not only do we recover the statistics for $a_{l}$’s, we have also recovered the feature learned by each branch.
4.2 Student-Teacher experiment on robust branch learning
We demonstrate this robust learning phenomenon and provide a first evidence of the equipartition conjecture with the student-teacher experiment setup introduced in the previous section. We use the contextual stochastic block model (CSBM) (Deshpande et al. (2018)) as the shared node input features, where the adjacency matrix is given by a stochastic block model with two blocks, and the node feature is generated with latent vectors corresponding to the two blocks (See B.1 for details).
Figure 2: Student readout norms squared as a function of network width. Left: Student branches break the GP symmetry as it goes to the narrow width limit. Middle, Right: Branch 0 and branch 1 readout norm squared for a range of $\sigma$ regularization values. Different regularization strengths will all converge to the same teacher readout norm values.
As shown in Figure 1(b), an extremely narrow student network learns the teacher’s branch readout norms very robustly, despite having a large variance; on the other hand, a much wider network fails to learn the teachers’ norms and approaches the GP limit, while having a smaller variance.
Symmetry Breaking and Convergence of Branches:
Figure 2 shows the student branch norms as a function of the network width. The symmetry breaking of branch norms from the GP limit to the narrow width limit accompanies the convergence to learning teacher’s norms at narrow width for different $\sigma$’s, supporting Eq. 21.
Narrow-to-Wide Width Transition:
Using the mean predictor and variance from the theory, we can determine the generalization of individual branches (Appendix A.2) as a function of network width $N$, as shown in Figure 3. At narrow width, we expect individual branch to learn the teacher’s branch output $y_{l}$ independently, causing the bias to increase with network width. This is observed for both branches, with a transition from the narrow-width regime to the GP regime. The regularization strength $\sigma^{2}$ controls the transition window, with larger $\sigma$’s leading to sharper transitions. This aligns with our analysis of the entropic and energetic contributions, where larger $\sigma$ amplifies the distinction between the two terms. In contrast, the variance decreases with network width for small $\sigma$’s, resulting in a trade-off between the contributions of bias and variance to overall generalization performance, as shown clearly with $\sigma=0.5$ in Figure 4.
Figure 3: Student network bias and variance for individual branch as a function of network width for student-teacher tasks.
Figure 4: Overall generalization performance as a function of network width for student-teacher tasks.
5 BPB-GNN on Cora
We also conduct experiments on the benchmark dataset Cora with binary node regression, and observe a similar narrow-to-wide width transition for the bias term. We analyze the generalization performance for both network width $N$ and the number of branches used $L$. As shown in Figure 5, the bias increases with network width, transitioning to the GP regime. Generally, more branches improve performance.
Figure 5: Cora generalization performance vs. network width $N$ and branch number $L$. The accuracy is computed by turning the mean predictor into a class label using its sign.
Convergence of Branch Importance at Narrow Width
An interesting aspect of the BPB-GNN network is that the branch norms seem to converge regardless of the hyperparameters $\sigma$ and $L$, reflecting the natural branch importance of the dataset and task.
Figure 6: Cora branch norms vs. network width $N$, $L=6$.
As shown in Figure 6, the BPB-GNN with $L=6$ branches robustly learns the readout norms at narrow width, verifying the equipartition conjecture. The last branch of the BPB-GNN network has a larger contribution, reflecting the presence of higher-order convolutions in the Cora dataset. From a kernel perspective, increasing branches better distinguish the nodes, as shown in Figure 7. This could explain the selective turn-off of intermediate branches and the increased contribution of the last branch.
Figure 7: Branch kernel $K_{l}$ visualizations the first 8 branches on Cora.
Figure 8: Branch Importance vs. L at narrow width.
Additionally, the first two branches are most robustly learned at narrow width, as shown in Figure 8, where the branch norms converge for the first two branches even for BPB-GNNs with different $L$. This suggests that the branch importance, as reflected by the norms learned at narrow width, indicates the contribution of the bare data and the first convolution layer.
6 Discussion
The findings presented in this paper reveal that BPB-GNNs exhibit unique characteristics in the narrow width limit. Unlike the infinite-width limit where neural networks behave as Gaussian Process (GP) with task-independent kernels, narrow-width BPB-GNNs undergo significant kernel renormalization. This renormalization leads to symmetry breaking among the branches, resulting in more robust and differentiated learning. Our experiments demonstrate that narrow-width BPB-GNNs can retain and, in some cases, improve generalization performance compared to their wider counterparts, particularly in bias-limited scenarios where the regularization effects dominate. Additionally, the observed independence of readout norms from architectural hyperparameters suggests that narrow-width BPB-GNNs can capture the intrinsic properties of the data more effectively. These insights suggest potential new strategies for optimizing neural network architectures in practical applications, challenging the traditional emphasis on increasing network width for better performance. Although our work is focused on GNN architectures, our findings can be transferred to the transformer family of architectures, as the transformer with frozen attention is analogous to a fully connected GNN (Vaswani et al. (2017); Veličković et al. (2017)), and a Bayesian theory with kernel renormalization can be similarly developed as in Tiberi et al. (2024).
Limitations:
We demonstrate that the bias term robustly decrease at narrow width; however, the variance term sometimes dominates and it is not clear BPB-GNN generalization is provably smaller at narrow width. Furthermore, our results rely on the Bayesian network setting, which might not directly transfer to understanding generalization of GNN trained with SGD, for which the loss landscape may be different.
7 Conclusion
In conclusion, this paper introduces and investigates the concept of narrow width limits in Bayesian Parallel Branching Graph Neural Networks. Contrary to the common belief that wider networks inherently generalize better, our results indicate that BPB-GNNs with significantly narrower widths can achieve better or competitive performance. This is attributed to effective symmetry breaking and kernel renormalization in the narrow-width limit, which lead to robust learning. Our theoretical analysis, supported by empirical evidence, establishes a new understanding of how network architecture influences learning outcomes. This work provides a novel perspective on neural network design and suggests further research into optimizing network structures for specific learning tasks, especially in scenarios where model simplicity and computational efficiency are crucial.
Acknowledgments and Disclosure of Funding
We acknowledge support of the Swartz Foundation, the Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University and the Gatsby Charitable Foundation. This material is partially based upon work supported by the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF-1231216. We have benefitted from helpful discussions with Alexander van Meegen, Lorenzo Tiberi and Qianyi Li.
Appendix A Details on Theory of BPB-GNN
A.1 Kernel renormalization
[…]
First, we linearize the energy in terms of $a_{l}$’s by introducing the auxiliary variables $t^{\mu},\mu=1,\dots,P$.
[…]
Now we can integrate out $a_{l}$’s as they are linearized and the partition function becomes
$Z=\int\,DWe^{-H(W)},$ (25)
with effective Hamiltonian
$H(W)=\frac{1}{2\sigma^{2}}\sum_{l=0}^{L-1}\mathrm{Tr}W_{l}^{T}W_{l}+\frac{1}{2}Y^{T}(K(W)+TI)^{-1}Y+\frac{1}{2}\log\det(K(W)+TI),$ (26)
where
$K(W)=\frac{1}{L}\sum_{l}\frac{\sigma^{2}}{N}(H_{l}(W_{l})H_{l}(W_{l})^{T})|_{P}$ (27)
is the $P\times P$ kernel matrix dependent on the observed $P$ nodes with node features $H_{l}=A^{l}XW_{l}$ and denote $|_{P}$ as the matrix restricting to the elements generated by the training nodes.
Now we perform the integration on $W_{l}$’s, and get a Fourier representation of $Z$ with $h_{l}$,$u_{l}$ as auxiliary variables after inserting $t$:
$Z$ $=\int\prod_{l=0}^{L}dh_{l}du_{l}dt\exp\Big{(}it^{T}Y-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}h_{l}-\frac{1}{2}t^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}t\Big{)}$
$=\int\prod_{l=0}^{L}dh_{l}du_{l}\exp\Big{(}-\sum_{l}\frac{N}{2}\log(1+h_{l})+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}h_{l}\frac{1}{2}Y^{T}\Big{(}\sum_{l}\frac{1}{L}u_{l}K_{l}+TI\Big{)}^{-1}Y\Big{)}$ (28)
where
$K_{l}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{lT}]|_{P}$ (29)
is the input kernel for branch $l$. Now as $N\to\infty$ and $\alpha=\frac{P}{N}$ fixed, we can perform the saddle point approximation and get the saddle points for $h_{l}$ as
$1+h_{l}=\frac{\sigma^{2}}{u_{l}}$ (30)
Plugging this back to the equation, we get
[…]
with the effective Hamiltonian
$H_{eff}=S+E,$ (32)
where we call $S$ the entropic term
$S=-\sum_{l}\frac{N}{2}\log u_{l}+\sum_{l}\frac{N}{2\sigma^{2}}u_{l}$ (33)
and $E$ the energetic term
$E=\frac{1}{2}Y^{T}(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)^{-1}Y+\frac{1}{2}\log\det(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)$ (34)
Therefore, after integrating out $W_{l}$, the effective kernel is given by
$K=\sum_{l}\frac{1}{L}u_{l}K_{l},$ (35)
where $K_{l}$ is
$K_{l}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{P}$ (36)
And the saddle point equations for $u_{l}$’s are determined by
$N(1-\frac{u_{l}}{\sigma^{2}})=-Y^{T}(K+TI)^{-1}\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y+\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}],$ (37)
where we call
$r_{l}=Y^{T}(K+TI)^{-1}\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y$ (38)
and
$\mathrm{Tr}_{l}=\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}]$ (39)
As we will show later, these represent the mean and variance of the readout norm squared respectively. In the $T=0$ case, the saddle point equation becomes
$N(1-\frac{u_{l}}{\sigma^{2}})=-Y^{T}K^{-1}\frac{u_{l}K_{l}}{L}K^{-1}Y+\mathrm{Tr}[K^{-1}\frac{u_{l}K_{l}}{L}]$ (40)
A.2 Predictor statistics and generalization
We can get the predictor statistics of each branch readout $y_{l}^{\nu}(G)$ on a new test node $\nu$ by considering the generating function:
$Z(\eta_{1},\dots,\eta_{L})=\int D\Theta\exp{\left\{-\frac{\beta}{2}\sum_{\mu}(f^{\mu}(G;\Theta)-y^{\mu})^{2}+\sum_{l}i\eta_{l}\frac{1}{\sqrt{NL}}\sum_{i}a^{(l)}_{i}h_{i}^{(l),\nu}(G,W_{l})-\frac{T}{2\sigma^{2}}\Theta^{T}\Theta\right\}}$ (41)
Therefore, by taking the derivative with respect to each $\eta_{l}$, we arrive at the statistics for $y_{l}(x)$ as:
$\langle y_{l}^{\nu}(G)\rangle=\partial_{i\eta_{l}}\log Z\big{|}_{\vec{\eta}=0}$ (42)
$\langle\delta y_{l,\nu}^{2}(G)\rangle=\partial^{2}_{i\eta_{l}}\log Z\big{|}_{\vec{\eta}=0}$ (43)
After integrating out the weights $\Theta$ layer by layer, we have:
$Z(\eta_{1},\dots,\eta_{L})$ $=\int\Pi_{l}du_{l}\,\exp{\left\{\sum_{l}\left(\frac{N}{2}\log u_{l}-\frac{N}{2\sigma^{2}}u_{l}\right)\right.}$
$+\frac{1}{2}(iY+\sum_{l}\frac{1}{L}\eta_{l}u_{l}k_{l}^{\nu})^{T}(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)^{-1}(iY+\sum_{l}\eta_{l}\frac{1}{L}u_{l}k_{l}^{\nu}$
$-\frac{1}{2}\log\det(\sum_{l}\frac{1}{L}u_{l}K_{l}+TI)-\frac{1}{2}\sum_{l}\eta_{l}^{2}\frac{1}{L}K^{\nu,\nu}_{l}\bigg{\}}.$ (44)
Here
$k_{l}^{\nu}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{(P,\nu)}$ (45)
is the $P\times 1$ column kernel matrix for test node $\nu$ and all training nodes, and
$K_{l}^{\nu,\nu}=\frac{\sigma^{2}}{N_{0}}[A^{l}XX^{T}A^{l}]|_{(\nu,\nu)}$ (46)
is the single matrix element for the test node. Therefore, eventually, we have:
$\langle y_{l}^{\nu}\rangle=\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}Y$ (47)
and
$\langle\delta y_{l,\nu}^{2}\rangle=\frac{u_{l}K_{l}^{\nu,\nu}}{L}-\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}\frac{u_{l}k_{l,\nu}}{L}$ (48)
The predictor statistics of the overall readout $y=\sum_{l}y_{l}$ is given by:
$\langle y^{\nu}(G)\rangle=\sum_{l}\frac{u_{l}k_{l,\nu}^{T}}{L}(K+TI)^{-1}Y=k_{\nu}^{T}(K+TI)^{-1}Y$ (49)
$\langle\delta y(G)_{\nu}^{2}\rangle=\sum_{l}u_{l}K_{l}^{\nu,\nu}-\sum_{l,l^{\prime}}u_{l}k_{l,\nu}^{T}(K+TI)^{-1}u_{l^{\prime}}k_{l^{\prime},\nu}=K_{\nu,\nu}-k_{\nu}^{T}(K+TI)^{-1}k_{\nu}$ (50)
A.3 Statistics of branch readout norms
From the partition function Eq.24, we can relate the mean of readout weights $a_{l}$ to the auxiliary variable $t$ by
$\langle a_{l}\rangle_{W}=-i\frac{\sigma^{2}}{\sqrt{N}}\Phi_{l}^{T}\langle t\rangle=-\frac{\sigma^{2}}{\sqrt{NL}}\Phi_{l}^{T}(K+TI)^{-1}Y,$ (51)
where $\Phi_{l}$ is the node feature matrix for the hidden layer nodes. We have
$\langle a_{l}^{T}\rangle\langle a_{l}\rangle=\sigma^{2}Y^{T}(K+TI)^{-1}\frac{u_{l}K_{l}}{L}(K+TI)^{-1}Y=r_{l}\sigma^{2}$ (52)
We can calculate the second-order statistics of $a_{l}$: the variance is
$\langle\delta a_{l}^{T}\delta a_{l}\rangle=\sigma^{2}\mathrm{Tr}(I+\frac{\sigma^{2}\beta}{NL}\Phi_{l}\Phi_{l}^{T})^{-1}\\ =\sigma^{2}(N-\mathrm{Tr}(K+TI)^{-1}\frac{u_{l}K_{l}}{L})=\sigma^{2}(N-\mathrm{Tr}_{l})$ (53)
Therefore,
$\langle a_{l}^{2}\rangle=\langle\delta a_{l}^{T}\delta a_{l}\rangle+\langle\delta a_{l}^{T}\delta a_{l}\rangle=N\sigma^{2}+\sigma^{2}r_{l}+1-\sigma^{2}\mathrm{Tr}_{l}=Nu_{l}$ (54)
Therefore, we have proved the main text claim that the order parameter $u_{l}$’s are really the mean squared readout norms of the branches.
Appendix B Experimental details
B.1 Student-teacher CSBM
For the student-teacher task, we use the contextual stochastic block model introduced by Deshpande et al. [2018]. The adjacency matrix is given by
$A_{ij}=\begin{cases}1&\text{with probability }p=c_{in}/n,\text{ if }i,j\leq n/2\\ 1&\text{with probability }p=c_{in}/n,\text{ if }i,j\geq n/2\\ 1&\text{with probability }q=c_{out}/n,\text{ otherwise}\end{cases}$ (55)
where
$c_{in,out}=d\pm\sqrt{d}\lambda$ (56)
$d$ is the average degree and $\lambda$ the homophily factor.
The feature vector $\vec{x}^{\mu}$ for a particular node $\mu$ is given by
$\vec{x}_{\mu}=\sqrt{\frac{\mu}{n}}y^{\mu}\vec{u}+\vec{\xi}_{\mu},$ (57)
where
$\vec{u}\sim\mathcal{N}(0,I_{N_{0}}),\vec{\xi}_{\mu}\sim\mathcal{N}(0,I_{N_{0}})$ (58)
In the experiment, we use $N_{0}=950$,$d=20$,$\lambda=4$ and $\mu=4$. The teacher network parameters are variance $\sigma_{t}^{2}=1$, width $N_{t}=1024$, branch norms variance $\beta_{0}^{2}=0.4$, $\beta_{1}^{2}=2$ for individual element of the readout vector $a_{l}$. Temperature $T=0.0005\sigma^{2}$ for each $\sigma$ value.
B.2 Cora
For the Cora dataset, we use a random split of the data into 21% as training set and 79% as test set. We also group the classes $(1,2,4)$ into one group and the rest for the other group for binary node regression. We use temperature $T=0.01$ for both theory and sampling as the sampling becomes more difficult for smaller temperature. This explains the discrepancy of the GP limit bias for different $\sigma$ values.
B.3 Hamiltonian Monte Carlo
The sampling in the paper are all done with Hamiltonian Monte Carlo simulations, which has fast convergence to the posterior distribution compared to Langevin dynamics. Due to memory constraint, we only sampled up to $N=1024$ hidden layer width for the student-teacher CSBM experiment and $N=64$ for the Cora experiment. Since we mainly aim to demonstrate the narrow width effect in this paper, this suffices the purpose.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See our experimental evidence
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See 6
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: See A
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: See B
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: Code not ready to be released.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: See B
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Hamiltonian Monte Carlo Experiments computationally expensive to measure error.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [No]
  39. Justification: Hamiltonian Monte Carlo experiments run over different compute budget.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We understand the code of ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We do not forsee imminent societal impact from the theoretical analysis.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [No]
  54. Justification: We do not forsee high risk release.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: No assets involved.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: No assets relesed.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: not crowdsourcing
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: no need for IRB approval
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.